# A Biologically Plausible Dense Associative Memory with Exponential Capacity

### v1

@@ Abstract
Krotov and Hopfield (2021) proposed a biologically plausible two-layer associative memory network with memory storage capacity exponential in the number of visible neurons. However, the capacity was only linear in the number of hidden neurons. This limitation arose from the choice of nonlinearity between the visible and hidden units, which enforced winner-takes-all dynamics in the hidden layer, thereby restricting each hidden unit to encode only a single memory. We overcome this limitation by introducing a novel associative memory network with a threshold nonlinearity that enables distributed representations. In contrast to winner-takes-all dynamics, where each hidden neuron is tied to an entire memory, our network allows hidden neurons to encode basic components shared across many memories. Consequently, complex patterns are represented through combinations of hidden neurons. These representations reduce redundancy and allow many correlated memories to be stored compositionally. Thus, we achieve much higher capacity: exponential in the number of hidden units, provided the number of visible units is sufficiently larger than the number of hidden neurons. Exponential capacity arises because all binary states of the hidden units can become stable memory patterns with an appropriately chosen threshold. Moreover, the distributed hidden representation, which has much lower dimensionality than the visible layer, preserves class-discriminative structure, supporting efficient nonlinear decoding. These results establish a new regime for associative memory, enabling high-capacity, robust, and scalable architectures consistent with biological constraints.
@@ 1 Introduction
Associative memory networks are a class of attractor models in which the system can recall stored memories from their incomplete or noisy versions via recurrent dynamics ([krotov2025modern]). In such models, memories are conceptualized as the stable fixed points of the network dynamics. The number of fixed points determines the storage capacity of the network, and significant efforts have been made to construct networks with sufficiently high storage capacity to explain human memory.
@@
The classical Hopfield network, a leading model for associative memory, has a storage capacity that scales linearly with the number of neurons in the network ([hopfield1982neural]). Dense associative memories, sometimes also referred to as modern Hopfield networks ([krotov2016dense]), are promising modifications of the classical Hopfield model. By incorporating higher-order interactions (e.g., interactions that are quadratic, rather than linear, in the input to a neuron), they achieve a storage capacity that scales super-linearly with the number of neurons. There are many possible choices for the energy function in this class of models. For instance, the power interaction vertex leads to the power-law scaling of the capacity ([baldi1987number, gardner1987multiconnected, abbott1987storage, horn1988capacities, chen1986high, krotov2016dense]). More sophisticated shapes of the energy function result in exponential storage capacity in the number of neurons, while maintaining large basins of attraction ([demircigil2017model, lucibello2024exponential]).
@@
The naive implementation of Dense Associative Memory models, however, relies on synaptic interactions that are challenging to implement broadly in biological circuits. In particular, these models require nonlinear interactions among synapses. While several biological mechanisms could in principle support restricted forms of higher-order interactions, such as astrocytic processes, dendritic computations, or distributed neurotransmitter effects, these remain limited in scope and dictate strong constraints on the possible shape of the energy landscape ([burnssimplicial, kozachkov2025neuron, kafrajbiologically]). The implementation of Dense Associative Memory introduced by [krotov2021large] does not suffer from these limitations, as it relies only on standard synaptic interactions. In this architecture, the visible neurons correspond to features of the patterns, while the hidden neurons serve as auxiliary computational elements that mediate complex interactions. Higher-order interactions among visible neurons emerge by selecting appropriate activation functions for the hidden neurons.
@@
Nevertheless, the two-layer implementation of Krotov and Hopfield has two key limitations. First, the storage capacity is at most linear in the number of hidden neurons ([krotov2021hierarchical, krotov2021large]). This is unsatisfactory from the perspective of information storage – one would like to store as much information as possible while utilizing only a small number of neurons. Second, at inference time, the network demonstrates a winner-takes-all behavior. This means that the asymptotic fixed point that the network converges to corresponds to a single hidden neuron being activated, while the rest of the hidden neurons are inactive. This behavior results in grandmother-like representations for hidden neurons, as opposed to distributed representations, which are more efficient at storing information.
@@
Our work tackles these two limitations. Specifically, we present a novel implementation of Dense Associative Memory that achieves exponential storage capacity in the number of hidden neurons. This is accomplished with a simple yet critical change: we use a threshold activation function that does not enforce winner-takes-all dynamics. The threshold activation enables distributed memory representations–multiple hidden neurons can be active for a memory, and each hidden neuron can participate in multiple memories. As a result, all possible binary patterns of hidden neuron states become stable fixed points, enabling the network to store exponentially many memories, including highly correlated ones. Beyond high capacity, the hidden layer of the network is low-dimensional compared to the visible layer, yet it produces structured representations that preserve class-discriminative information, with memories sharing components represented close together in the hidden activity space. We establish this result through both theoretical analysis and numerical simulations, and show that the resulting fixed points also possess large basins of attraction.
@@
Our model is closely related to the framework recently proposed by [chandra2025episodic], which combines multiple Dense Associative Memory modules to produce a distributed code for the visible neurons. Each module performs a winner-takes-all operation similar to [krotov2021large], so only a single hidden neuron is active per module. By combining several modules, though, they achieve exponential storage capacity. However, we show that multiple modules are unnecessary: exponential capacity can be achieved with a single module, provided the activation function is chosen appropriately.
@@
Beyond its biological motivation, our work also connects to a growing body of research on Dense Associative Memories in machine learning. Notably, it has been shown that Dense Associative Memory closely corresponds to the attention mechanism in transformer architectures [ramsauerhopfield, hoover2023energy], offering a principled framework for viewing the transformer’s attention and feedforward computations as steps in a global energy minimization process. Complementary research has demonstrated that generative diffusion models, widely used in high-quality image generation, also exhibit associative memory behavior [hoover2023memory, ambrogioni2024search, pham2025memorization]. Further studies have expanded the model’s functionality: for instance, [chaudhry2023long] examined its ability to store and retrieve long sequences; [burnssimplicial] introduced higher-order simplicial interactions; and [dohmatob2023different, hoover2025dense] proposed alternative energy functions that also support exponential storage capacity. Our results contribute to this line of work by showing how exponential storage capacity can be achieved within a biologically plausible two-layer framework, thereby bridging theoretical neuroscience with modern machine learning architectures.
@@ 2 Model
In this section, we present our model and demonstrate that its storage capacity scales exponentially with the number of hidden neurons, meaning that all possible binary patterns of hidden neurons are stable fixed points.
We first define the dynamics of the system as follows:
@@
where $\Theta(\cdot)$ is the standard Heaviside step function:
@@
Synaptic connections between visible neuron $i$ and hidden neuron $\mu$ are reciprocal and randomly drawn from a standard normal distribution:
@@
The scaling factors in front of the sums are chosen purely for convenience, as they simplify subsequent expressions. Additionally, for simplicity, our theoretical analysis and experiments assume a Heaviside step function; however, Appendix F shows that this assumption is not strictly required.
@@ 2.1 Storage Capacity
Equation (5), with the weight matrix given in Eq. (6), is very close to the classical Hopfield model; the only difference is that in the classical model, the ${{\xi}}_{\mu i}$ are binary, whereas in our model they’re Gaussian. However, the classical Hopfield model works in the regime $N_{v}<N_{h}$, with memory storage possible only if $N_{v}<0.138N_{h}$([amit1985storing]). Here, though, we’ll consider a very different regime: $N_{v}\gg N_{h}$. In this limit, $J_{\mu\nu}$ approaches the identity matrix ([marchenko1967distribution]), which completely decouples the hidden neurons. Assuming the threshold, $\theta$, is chosen correctly, this leads immediately to exponential storage capacity.
@@
$J_{\mu\nu}=\delta_{\mu\nu}+\frac{\zeta_{\mu\nu}}{\sqrt{N_{v}}}$ (7)
where the $\zeta_{\mu\nu}$ are independent, zero-mean, unit-variance Gaussian random variable,
$\zeta_{\mu\nu}\sim\mathcal{N}(0,1)\,$, (8)
and here and in what follows $\delta_{\mu\nu}$ is the Kronecker delta. Thus, Eq. (5) may be written
@@
$|q_{\mu}|\equiv\left|\frac{1}{\sqrt{N_{v}}}\sum_{\nu=1}^{N_{h}}\zeta_{\mu\nu}s_{\nu}\right|\sim\sqrt{\frac{1}{N_{v}}\sum_{\nu=1}^{N_{h}}s_{\nu}^{2}}\leq\sqrt{\frac{N_{h}}{N_{v}}}$ (10)
@@
If we set $\theta=1/2$, in the limit $N_{v}\gg N_{h}$ Eq. (9) typically has two solutions: one at $s_{\mu}=0$ and one at $s_{\mu}=1$. In fact, the probability that there is only one solution is the probability that $(2s_{\mu}-1)q_{\mu}<-1$; in Appendix D (see in particular Eq. (34)) we show that
$P_{\text{no flip}}=\Phi\left(\sqrt{\frac{N_{v}}{4N_{a}}}\right)^{N_{h}}\,$ (11)
where $N_{a}$ is the number of active hidden neurons. Thus, even when $N_{v}$ is only about ten times larger than $N_{h}$, and the threshold is not exactly 1/2, there are approximately $2^{N_{h}}$ fixed points.
@@
Since $J_{\mu\nu}$ is approximately the identity matrix, we see that at equilibrium $h_{\mu}$ is close to either 0 or 1. Thus, because our nonlinearity is a step function, its derivative vanishes at equilibrium, guaranteeing the stability of the fixed points (Appendix A.2). Consequently, when we solve Eq. (1), we expect to see $2^{N_{h}}$ stable fixed points in the regime $N_{v}\gg N_{h}$. This prediction is consistent with numerical simulations, as can be seen in Figure 1a.
@@ 2.2 Basins of Attraction
where $h_{\mu,\text{target}}=1$ if neuron $\mu$ encodes the target memory, and 0 otherwise (motivated by the fact that $h_{\mu}$ is close to either 0 or 1 at the fixed points; see Eq. (12)).
@@
Using Eq. (7), we see that the first term is $\Theta(h_{\mu,\text{target}}-\theta)+\mathcal{O}(\sqrt{N_{h}/N_{v}})$. And the second term scales as $\sigma_{v}\sqrt{N_{h}/N_{v}}$ where $\sigma_{v}^{2}$ is the variance of the noise. Thus, so long as
$\text{Var}[\epsilon]\ll\frac{N_{v}}{N_{h}}\,$, (15)
$h_{\mu}(t)$ will be close to its target value when $t\ll\tau_{v}$. Since $v_{i}(t)$ is close to its target value at that time, it will stay close, and asymptotically the target pattern will be recovered. Given that $N_{v}\gg N_{h}$, $\epsilon_{i}^{v}$ can be very large without affecting recall. Thus, the basin of attraction is very large (see Figure 1b and Appendix D).
@@ 2.3 Biological Plausibility
Compared to [krotov2021large], our model exhibits greater biological plausibility in several respects.
The activation function used here is local and keeps neuron activity within a biologically realistic range. In contrast, in [krotov2021large], Model A is not biologically plausible because the hidden neuron activities can grow to unrealistically large values as a consequence of the power-law activation, which does not reflect realistic neural firing. Models B and C, on the other hand, rely on non-local activation functions, softmax and spherical normalization, respectively, which are biologically implausible unless additional mechanisms are assumed.(see Appendix C for more details).
Although our theoretical analysis focuses on symmetric weights and a global threshold for all neurons for simplicity, these assumptions are not restrictions of the model. Experimentally, we show that networks with asymmetric weights and heterogeneous neuron-specific thresholds also achieve stable recall. Allowing asymmetric weights is important because exact symmetry is rarely observed in biological neural circuits, yet memory networks can remain robust even without it. Similarly, heterogeneous thresholds capture the variability in neuron excitability across real neurons and demonstrate that stable memory dynamics do not require finely tuned, uniform parameters. Together, these features indicate that our model better reflects realistic neural mechanisms while retaining associative memory functionality. Figure 9 in Appendix E shows representative recall examples for networks with asymmetric weights and heterogeneous thresholds that stored MNIST and CIFAR-10 images, respectively, using a learning rule similar to that discussed in Subsection 3.1, with the key differences being the absence of a symmetry restriction on the weights and the allowance of heterogeneous thresholds.
@@ 3.1 Learning Rule
In the steady-state, visible activity in Eq. (1) can be expressed as
$\mathbf{v}=\frac{1}{\sqrt{N_{h}}}\sum_{\mu=1}^{N_{h}}\boldsymbol{\xi}_{\mu}s_{\mu},$ (16)
where $\boldsymbol{\xi}_{\mu}$ is the $\mu$-th column of $\xi\in\mathbb{R}^{N_{v}\times N_{h}}$, i.e., $(\boldsymbol{\xi}_{\mu})_{i}=\xi_{i\mu}$. If only hidden neuron $\mu$ is active, the visible state equals $\boldsymbol{\xi}_{\mu}$. A visible memory is thus called basic if it corresponds to a single active hidden neuron, and complex if it is formed by the activation of multiple hidden neurons, i.e. a composition of several basic memories.
@@
$(\xi,\theta)=\arg\min_{\xi,\theta}\sum_{m=1}^{M}\Big\|\mathbf{v}_{m}-\frac{1}{\sqrt{N_{h}}}\sum_{\mu=1}^{N_{h}}\boldsymbol{\xi}_{\mu}\,\Theta(\frac{\sqrt{N_{h}}}{N_{v}}\boldsymbol{\xi}_{\mu}^{\top}\mathbf{v}_{m}-\theta)\Big\|^{2},$ (17)
where $s_{\mu}$ is replaced by its target steady-state value. This learning rule is identical to the one proposed in [radhakrishnan2020overparameterized]. We used Xavier initialization for the weights and approximated the threshold function $\Theta$ with a sharp sigmoid to allow gradient-based training.

### v2

@@ Abstract
Krotov and Hopfield (2021) proposed a biologically plausible two-layer associative memory network with memory storage capacity exponential in the number of visible neurons. However, the capacity was only linear in the number of hidden neurons. This limitation arose from the choice of nonlinearity between the visible and hidden units, which enforced winner-take-all dynamics in the hidden layer, thereby restricting each hidden unit to encode only a single memory. We overcome this limitation by introducing a novel associative memory network with a threshold nonlinearity that enables distributed representations. In contrast to winner-take-all dynamics, where each hidden neuron is tied to an entire memory, our network allows hidden neurons to encode basic components shared across many memories. Consequently, complex patterns are represented through combinations of hidden neurons. These representations reduce redundancy and allow many correlated memories to be stored compositionally. Thus, we achieve much higher capacity: exponential in the number of hidden units, provided the number of visible units is sufficiently large relative to the number of hidden units. Exponential capacity arises because all binary states of the hidden units can become stable memory patterns. Moreover, the distributed hidden representation, which has much lower dimensionality than the visible layer, preserves class-discriminative structure, supporting efficient nonlinear decoding. These results establish a new regime for associative memory, enabling high-capacity, robust, and scalable architectures consistent with biological constraints.
@@ 1 Introduction
Associative memory networks are a class of attractor models in which the system can recall stored memories from their incomplete or noisy versions via recurrent dynamics (Krotov et al. (2025)). In such models, memories are conceptualized as the stable fixed points of the network dynamics. The number of fixed points determines the storage capacity of the network, and significant effort has been made to construct networks with sufficiently high storage capacity to explain human memory.
@@
The classical Hopfield network, a leading model for associative memory, has a storage capacity that scales linearly with the number of neurons in the network (Hopfield (1982)). Dense associative memories, sometimes also referred to as modern Hopfield networks (Krotov and Hopfield (2016)), are promising modifications of the classical Hopfield model. By incorporating higher-order interactions (e.g., interactions that are quadratic, rather than linear, in the input to a neuron), they achieve a storage capacity that scales super-linearly with the number of neurons. There are many possible choices for the energy function in this class of models. For instance, the power interaction vertex leads to the power-law scaling of the capacity (Baldi and Venkatesh (1987); Gardner (1987); Abbott and Arian (1987); Horn and Usher (1988); Chen et al. (1986); Krotov and Hopfield (2016)). More sophisticated shapes of the energy function result in exponential storage capacity in the number of neurons, while maintaining large basins of attraction (Demircigil et al. (2017); Lucibello and Mézard (2024)).
@@
The naive implementation of Dense Associative Memory models, however, relies on synaptic interactions that are challenging to implement in biological circuits. In particular, these models require nonlinear interactions among synapses. While several biological mechanisms could in principle support restricted forms of higher-order interactions, such as astrocytic processes, dendritic computations, or distributed neurotransmitter effects, these remain limited in scope and dictate strong constraints on the possible shape of the energy function (Burns and Fukai (2023); Kozachkov et al. (2025); Kafraj et al. (2025)). The implementation of Dense Associative Memory introduced by Krotov and Hopfield (2021) does not suffer from these limitations, as it relies only on standard synaptic interactions. In this architecture, the visible neurons correspond to features of the patterns, while the hidden neurons serve as auxiliary computational elements that mediate complex interactions. Higher-order interactions among visible neurons emerge by selecting appropriate activation functions for the hidden neurons.
@@
Nevertheless, the implementation of Krotov and Hopfield has two key limitations. First, the storage capacity is at most linear in the number of hidden neurons (Krotov (2021); Krotov and Hopfield (2021)). This is unsatisfactory from the perspective of information storage – one would like to store as much information as possible while utilizing only a small number of neurons. Second, at inference time, the network demonstrates a winner-take-all behavior. This means that the asymptotic fixed point that the network converges to corresponds to a single hidden neuron being activated, while the rest of the hidden neurons are inactive. This behavior results in grandmother-like representations for hidden neurons, as opposed to distributed representations, which are more efficient at storing information.
@@
Our work tackles these two limitations. Specifically, we present a novel implementation of Dense Associative Memory that achieves exponential storage capacity in the number of hidden neurons. This is accomplished with a simple yet critical change: we use a threshold activation function that does not enforce winner-take-all dynamics. The threshold activation enables distributed memory representations–multiple hidden neurons can be active for a memory, and each hidden neuron can participate in multiple memories. As a result, all possible binary patterns of hidden neuron states become stable fixed points, enabling the network to store exponentially many memories, including highly correlated ones. Beyond high capacity, the hidden layer of the network is low-dimensional compared to the visible layer, yet it produces structured representations that preserve class-discriminative information, with memories sharing components represented close together in the hidden activity space. We establish this result through both theoretical analysis and numerical simulations, and show that the resulting fixed points also possess large basins of attraction.
@@
Our model is closely related to the framework recently proposed by Chandra et al. (2025), which combines multiple Dense Associative Memory modules to produce a distributed code for the visible neurons. In that work, each module performs a winner-take-all operation similar to Krotov and Hopfield (2021), so only a single hidden neuron is active per module. By combining several modules, though, they achieve exponential storage capacity. However, we show that multiple modules are unnecessary: exponential capacity can be achieved with a single module, provided the activation function is chosen appropriately.
@@
Beyond its biological motivation, our work also connects to a growing body of research on Dense Associative Memories in machine learning. Notably, it has been shown that Dense Associative Memory closely corresponds to the attention mechanism in transformer architectures (Ramsauer et al. (2021); Hoover et al. (2023a)), offering a principled framework for viewing the transformer’s attention and feedforward computations as steps in a global energy minimization process. Complementary research has demonstrated that generative diffusion models, widely used in high-quality image generation, also exhibit associative memory behavior (Hoover et al. (2023b); Ambrogioni (2024); Pham et al. (2025)). Further studies have expanded the model’s functionality: for instance, Chaudhry et al. (2023) examined its ability to store and retrieve long sequences; Dohmatob (2023) and Hoover et al. (2025) proposed alternative energy functions that also support exponential storage capacity; and Burns and Fukai (2023) introduced higher-order simplicial interactions. Our results contribute to this line of work by showing how exponential storage capacity can be achieved within a biologically plausible two-layer framework, thereby bridging theoretical neuroscience with modern machine learning architectures.
@@ 2 Model
In this section we present our model and demonstrate that its storage capacity scales exponentially with the number of hidden neurons, meaning that all possible binary patterns of hidden neurons are stable fixed points.
We first define the dynamics of the system as follows,
@@
where $\Theta(\cdot)$ is the standard Heaviside step function,
@@
Synaptic connections between visible neuron $i$ and hidden neuron $\mu$ are symmetric and drawn randomly from a standard normal distribution,
@@
The scaling factors in front of the sums in Eq. (1) are chosen purely for convenience, as they simplify subsequent expressions. Additionally, for simplicity, our theoretical analysis and experiments assume a Heaviside step function; however, Appendix F shows that this assumption is not required.
@@ 2.1 Storage Capacity
Equation (5), with the weight matrix given in Eq. (6), is very close to the classical Hopfield model; the only difference is that in the classical model, the ${{\xi}}_{\mu i}$ are binary, whereas in our model they’re Gaussian. However, the classical Hopfield model works in the regime $N_{v}<N_{h}$, with memory storage possible only if $N_{v}<0.138N_{h}$(Amit et al. (1985)). Here, though, we’ll consider a very different regime: $N_{v}\gg N_{h}$. In this limit, $J_{\mu\nu}$ approaches the identity matrix (Marchenko and Pastur (1967)), which decouples the hidden neurons. Assuming the threshold, $\theta$, is chosen correctly, this leads immediately to exponential storage capacity.
@@
$J_{\mu\nu}$ $=\delta_{\mu\nu}+\frac{\zeta_{\mu\nu}}{\sqrt{N_{v}}}$ (7a)
$\zeta_{\mu\nu}$ $\sim\mathcal{N}(0,1+\delta_{\mu\nu})\,.$ (7b)
Here and in what follows $\delta_{\mu\nu}$ is the Kronecker delta: it’s 1 if $\mu=\nu$ and 0 otherwise. Thus, Eq. (5) can be written as
@@
$|q_{\mu}|\equiv\left|\frac{1}{\sqrt{N_{v}}}\sum_{\nu=1}^{N_{h}}\zeta_{\mu\nu}s_{\nu}\right|\sim\sqrt{\frac{1}{N_{v}}\sum_{\nu=1}^{N_{h}}s_{\nu}^{2}(1+\delta_{\mu\nu})}\leq\sqrt{\frac{N_{h}+1}{N_{v}}}$ (9)
@@
If we set $\theta=1/2$, in the limit $N_{v}\gg N_{h}$ Eq. (8) typically has two solutions: one at $s_{\mu}=0$ and one at $s_{\mu}=1$. However, if $q_{\mu}>1/2$ then the only solution is $s_{\mu}=1$, and if $q_{\mu}<-1/2$ then the only solution is $s_{\mu}=0$. In both cases there can be a “bit flip”, where $s_{\mu}$ is forced to the wrong value. In Appendix D (see in particular Eqs. (40) and (36), but with $\sigma_{v}=0$) we show that
$P_{\text{no bit flips}}\geq 1-N_{h}\sqrt{\frac{N_{h}+1}{N_{v}}}\ \frac{e^{-\frac{N_{v}}{8(N_{h}+1)}}}{\sqrt{\pi/2}}\,.$ (10)
This bound approaches 1 exponentially rapidly as the number of visible units, $N_{v}$, increases.
@@
Since $J_{\mu\nu}$ is approximately the identity matrix, we see that at equilibrium $h_{\mu}$ is close to either 0 or 1. Because our nonlinearity is a step function, its derivative vanishes at equilibrium, guaranteeing the stability of the fixed points (Appendix A.2). Consequently, when we solve Eq. (1), we expect to see $2^{N_{h}}$ stable fixed points in the regime $N_{v}\gg N_{h}$. This prediction is consistent with numerical simulations, as can be seen in Figure 1a.
@@ 2.2 Basins of Attraction
where $h_{\mu,\text{target}}=1$ if neuron $\mu$ encodes the target memory, and 0 otherwise (motivated by the fact that $h_{\mu}$ is close to either 0 or 1 at the fixed points; see Eq. (11)) and the noise, $\epsilon_{i}^{v}$ is independent and normally distributed with variance $\sigma_{v}^{2}$,
$\epsilon_{i}^{v}\sim\mathcal{N}(0,\sigma_{v}^{2})\,.$ (13)
@@
Using Eq. (7a), we see that the first term is $\Theta(h_{\mu,\text{target}}-\theta)+\mathcal{O}(\sqrt{N_{h}/N_{v}})$. And the second term scales as $\sigma_{v}\sqrt{N_{h}/N_{v}}$. Thus, so long as
$\sigma_{v}^{2}\ll\frac{N_{v}}{N_{h}}\,$, (15)
$h_{\mu}(t)$ will be close to its target value when $t\ll\tau_{v}$. Since $v_{i}(t)$ is close to its target value at that time, it will stay close, and asymptotically the target pattern will be recovered. Given that $N_{v}\gg N_{h}$, the added noise can be very large without affecting recall. Thus, the basin of attraction is very large (see Figure 1b and Appendix D).
@@ 2.3 Biological Plausibility
Compared to Krotov and Hopfield (2021), our model exhibits greater biological plausibility in several respects.
The activation function used here is local and keeps neuron activity within a biologically realistic range. In contrast, in Krotov and Hopfield (2021), Model A is not biologically plausible because the hidden neuron activities can grow to unrealistically large values as a consequence of the power-law activation, which does not reflect realistic neural firing. Models B and C, on the other hand, rely on non-local activation functions, softmax and spherical normalization, respectively, which are biologically implausible unless additional mechanisms are assumed.(see Appendix C for more details).
Although our theoretical analysis focuses on symmetric weights and a global threshold for all neurons, these assumptions are not restrictions of the model. Experimentally, we show that networks with asymmetric weights and heterogeneous thresholds also achieve stable recall. Allowing asymmetric weights is important because exact symmetry is rarely observed in biological neural circuits. Similarly, heterogeneous thresholds capture the variability in neuron excitability across real neurons, and demonstrate that stable memory dynamics do not require finely tuned, uniform parameters. Together, these features indicate that our model better reflects realistic neural mechanisms while retaining associative memory functionality. Figure 9 in Appendix E shows representative recall examples for networks with asymmetric weights and heterogeneous thresholds that stored MNIST and CIFAR-10 images, respectively, using a learning rule similar to that discussed in Section 3.1.
@@ 3.1 Learning Rule
In steady-state, visible activity in Eq. (1) can be expressed as
$\mathbf{v}=\frac{1}{\sqrt{N_{h}}}\sum_{\mu=1}^{N_{h}}\bm{\xi}_{\mu}s_{\mu},$ (16)
where $\bm{\xi}_{\mu}$ is the $\mu$-th column of $\xi\in\mathbb{R}^{N_{v}\times N_{h}}$, i.e., $(\bm{\xi}_{\mu})_{i}=\xi_{i\mu}$. If only hidden neuron $\mu$ is active, the visible state equals $\bm{\xi}_{\mu}$. A visible memory is called basic if it corresponds to a single active hidden neuron, and complex if it is formed by the activation of multiple hidden neurons, i.e. a composition of several basic memories.
@@
$(\xi,\theta)=\arg\min_{\xi,\theta}\sum_{m=1}^{M}\left\|\mathbf{v}_{m}-\frac{1}{\sqrt{N_{h}}}\sum_{\mu=1}^{N_{h}}\bm{\xi}_{\mu}\,\Theta\left(\frac{\sqrt{N_{h}}}{N_{v}}\bm{\xi}_{\mu}^{\top}\mathbf{v}_{m}-\theta\right)\right\|^{2},$ (17)
where we have replaced $s_{\mu}$ with its target steady-state value. This learning rule is identical to the one proposed in Radhakrishnan et al. (2020). We used Xavier initialization for the weights and approximated the threshold function $\Theta$ with a sharp sigmoid to allow gradient-based training.